OpenStudy (linstead_cpd_freak):

A. 48 B. 36 C. 9 D. 16

OpenStudy (linstead_cpd_freak):

What is the solution of 3/4x = 12

jimthompson5910 (jim_thompson5910):

The equation is this \[\Large \frac{3}{4}x = 12\] right?

OpenStudy (linstead_cpd_freak):

yes

jimthompson5910 (jim_thompson5910):

The first step is to multiply both sides by 4 \[\Large \frac{3}{4}x = 12\] \[\Large 4*\frac{3}{4}x = 4*12\] \[\Large 3x = 48\] notice how the 4's on the left side cancel out which eliminates the fraction

OpenStudy (mathmale):

Unfortunately, 3/4x = 12 is ambiguous. One wonders whether you mean (3/4)x = 12 or 3/(4x) = 12. Please use parentheses to impart added clarity.

OpenStudy (mathmale):

Muliplying both sides by 4 clears your equation of fractions, which is helpful.

OpenStudy (linstead_cpd_freak):

So D right?

jimthompson5910 (jim_thompson5910):

yep, divide both sides by 3 to get x = 16

OpenStudy (linstead_cpd_freak):

Okay thank you!

jimthompson5910 (jim_thompson5910):

no problem